Solve and check each equation for x
x/x-2 - x+1/x = 8/x^2-2x
Thank You!!!

Answer by stanbon(48579)

x/(x-2) - (x+1)/x = 8/(x^2-2x)
LCM = x(x-2)
Multiply both sides by the LCM to get:
x^2 - (x+1)(x-2)=8
x^2-x^2+x+2=8
x=6
Cheers
Stan H.